OpenStudy (anonymous):

mark can walk to the top of a mountain and return in a total time of 12 hours and walked a total of 32 miles. His speed downhill is 2mph faster than his uphill. what is his speed downhill?

OpenStudy (phi):

how far did you get? any equations?

OpenStudy (anonymous):

i posted two questions are you asking about my first question

OpenStudy (phi):

asking about this one

OpenStudy (anonymous):

this is my first time on this sight so please be patient with me I think your asking about my second post?

OpenStudy (anonymous):

are you still there?

OpenStudy (phi):

Here are some of the equations: total time of 12 hours means Tu+Td= 12 (Tu time up, Td time down) Du= Dd = 32/2 (same distance up and down) speed downhill is 2mph faster than his uphill mean Vd= Vu-2 or Vu= Vd -2 (speed up is slower than down. makes sense)

OpenStudy (phi):

* Vd= Vu-2 ==> Vd= Vu+2 (Vd is faster)

OpenStudy (phi):

can you make any sense of this?

OpenStudy (anonymous):

ok i will set up my equation and try to solve;i'm not to good with word problems. thanks

OpenStudy (anonymous):

no, i'm totally lost

OpenStudy (anonymous):

but i will try

OpenStudy (phi):

what is his speed downhill? we want to find Vd in time distance speed problems immediately thing speed*time= distance

OpenStudy (phi):

*think

OpenStudy (phi):

you should try to find 2 equations. One for uphill and one for downhill

OpenStudy (anonymous):

ok

OpenStudy (phi):

if you think you have them, post them

OpenStudy (anonymous):

ok

OpenStudy (phi):

how about: Vd*Td= 16 (speed * time = distance) and Vu * Tu = 16

OpenStudy (anonymous):

did you get 16 by dividing 32 by 2

OpenStudy (anonymous):

the equation i set up was 12+2x=32 , is this wrong

OpenStudy (phi):

to the top of a mountain and return. The distance up must be the same as the distance down. total is 32 miles, so 16 up and 16 down. as for 12+2x=32, no, this is not how to think of the problem. Stick with speed * time= distance (and label what your variables stand for)

OpenStudy (anonymous):

ok i will re- work it

OpenStudy (phi):

You start with this pair Vd * Td= 16 Vu * Tu = 16 we want Vd (speed going down), so we want to get rid of Vu and Tu in the second equation Do you know how to do that, using these equations: Tu+Td= 12 or Tu= 12-Td Vu= Vd -2

OpenStudy (anonymous):

yes

OpenStudy (phi):

what do you get?

OpenStudy (anonymous):

i'm sorry, i'm totally lost and now i'm frustrated

OpenStudy (anonymous):

i will try later thanks

OpenStudy (phi):

start with Vu * Tu = 16 we also know Vu= Vd -2 and Tu= 12-Td so replace Vu in the equation with its equal (Vd-2). same for Tu. Can you do that?

OpenStudy (anonymous):

thanks, for your help, i'm taking a break i will go over the tips you gave me and try again in a few